OpenStudy (anonymous):

Simplify (2b2)4. Answer 2b6 8b6 8b8 16b8

OpenStudy (anonymous):

as in b squared?

OpenStudy (anonymous):

can you draw it for me

OpenStudy (anonymous):

Could you show us what you've done so far please?

OpenStudy (anonymous):

i dont know how to do this

OpenStudy (anonymous):

draw it out please

OpenStudy (anonymous):

Okay: Using your knowledge about exponents: \[(ab^n)^4=a^4b^{4n}\] What would that become? \[(2b^2)^4=?\]